OpenStudy (anonymous):

A kite is currently flying at an altitude of 19 meters above the ground. If the angle of elevation from the ground to the kite is 33 degrees find the length of the kite string to the nearest meter. a.35 meters b. 23 meters c. 10 meters d.29 meters

OpenStudy (anonymous):

 like that?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

ok skip that, use sine

OpenStudy (anonymous):

is it clear that \[\sin(33^\circ)=\frac{19}{x}\]?

OpenStudy (anonymous):

i got 35

OpenStudy (anonymous):

did you compute \[\frac{19}{\sin(33^\circ)}\]?

OpenStudy (anonymous):

yeah looks like \(34.88\) or \(35\) rounded